One of the two 10G Links into our Huntingdon Data Centre – running from Huntingdon to Manchester is currently down. This is primarily a backup link.
However this does mean that both Huntingdon and Manchester have reduced resilience currently and are ‘at-risk’
The fault has been raised to the carrier and they have already commenced work to restore service
We will update this post as we know more
Update from our suppliers:
Our engineers have arrived on site and performing site checks. We have scheduled for splicers to be dispatched and are awaiting their ETA.
Next Update Due: 01:00 26/06
Update from our suppliers
Our engineers have found a fibre event 6km from the Leicester exchange. Engineers are making they way to the area where there is a confirmed fibre break.
Our field teams have located damage on Aylestone road at the junction of Putney Rd West. Road works have cut and dug up multiple cables causing significant damage to cables and joints. It will take longer than expected to assess the full damage before a fix plan can be drawn out.
1 x 160f
4 x 24f
4 x 144f
2 x 8f
1 x 96f
2 out of 4 sub ducts damaged
Road works. Cables chewed up by a rotary cutting machine. Severe damage has been caused to cables.
construction have completed installation of x 2 new ducts
Cable gang are onsite.
replacement cable has been ordered.
Pull in replacement cables when it arrives into the adjacent joints and resplice to restore.
Replacement Cable will start arrviing between 6:45 – 7:00
Once it has arrived the cable gang can start the process of pulling it in.
Our suppliers are continuing to replace the damaged fibres into the new ducts. Once this is completed they will splice the new fibre
The Civil work needed to install this fibre is complete and the replacement fibre is in the ground. Work to splice the new fibre is ongoing. Given there are over 600 fibre pairs affected this will take several hours to complete.
We will continue to update as we have significant information
[Update 11:00 27/6]
All replacement fibres are in place and splicing has completed
However we are still seeing business impact and have sent ground teams to investigate a Transmode link that has remained down.
Once the investigating engineers return with more information we will be in touch to advise of our next actions.
[Update 12:00 27/6]
We currently have Field engineer red light testing the span of the Transmode link.
Once we have further information we will update you again.
[Update 17:00 27/6]
The Link has been up since 13:00 today – there has been some disturbance however this has been stable since 16:30 today. The suppliers have stated that all work has now been completed on this complex fault. We will monitor the link but we believe this is now stable and we have full resilience to our Huntingdon & Manchester loocations